# Gauges and Accelerated Optimization over Smooth and/or Strongly Convex Sets

### v3

@@ 3.1 Examples
$\mu$ $=\inf_{\|\bar{y}-b\|_{2}=1}\frac{1}{(\bar{y}-b)^{T}\bar{y}+\|\bar{y}\|^{2}_{2}}=\frac{1}{(1+\|b\|_{2})(2+\|b\|_{2})}\ ,$
$L$ $=\sup_{\|\bar{y}-b\|_{2}=1}\frac{(\bar{y}-b)^{T}\bar{y}+\|\bar{y}\|_{2}^{2}}{((\bar{y}-b)^{T}\bar{y})^{3}}=\frac{2-\|b\|_{2}}{(1-\|b\|_{2})^{2}}\ .$
Then following Lemma 15, $\frac{1}{2}\gamma_{E_{2}}^{2}$ is $\frac{\lambda_{\min}(A^{T}A)}{(1+\|b\|_{2})(2+\|b\|_{2})}$-strongly convex and $\frac{\lambda_{\max}(A^{T}A)(2-\|b\|_{2})}{(1-\|b\|_{2})^{2}}$-smooth.
@@ 3.3 Proofs of Theorems Characterizing Structured Gauges
Both strongly convex sets and smooth sets are defined in terms of balls built by a boundary point and corresponding normal vector. This perspective is critical for the proofs of our main theorems. The following lemmas characterize the gauges of these approximating balls. Their proofs are deferred to Appendix A.2.6 and A.2.7. Note that below, we consider sets $B(\bar{y}-r\zeta,r)$, which may not contain the origin. As a result, it is important that we defined the gauge as $\inf\{\lambda>0\mid x\in\lambda S\}$, which may no longer equal $\sup\{\lambda>0\mid x\not\in\lambda S\}$.
@@ 3.3.1 Proof of Theorem 11
$\partial\left(\frac{1}{2}\gamma^{2}_{B}\right)(y)=\gamma_{B}(y)\partial\gamma_{B}(y)=\left\{\frac{\gamma_{B}(y)\zeta}{\zeta^{T}\bar{y}}\right\}\ .$
@@
Since $S$ is strongly convex w.r.t. $(\bar{y},\zeta)$, then $S\cap B(\bar{y},\bar{\eta})\subseteq B\left(\bar{y}-\frac{1}{\alpha}\zeta,\frac{1}{\alpha}\right)\cap B(\bar{y},\bar{\eta})$ for some $\bar{\eta}$. Hence for $z$ near $y$, $\gamma_{S}(z)\geq\gamma_{B}(z)$. Applying this bound and Lemma 19 (a) and (b), we conclude for any $\tilde{\mu}<\mu$ in some neighborhood of $y$, all $z$ have
$\frac{1}{2}\gamma_{S}^{2}(z)\geq\frac{1}{2}\gamma_{B}^{2}(z)\geq\frac{1}{2}\gamma_{S}^{2}(y)+g^{T}(z-y)+\frac{\tilde{\mu}}{2}\|z-y\|^{2}\ .$ (19)
@@
If $\gamma_{S}(y)=0$ and $y=0$, then $\partial(\frac{1}{2}\gamma_{S}^{2})(y)=\{0\}$. For any point $z$, Lemma 5 ensures $\frac{1}{2}\gamma_{S}^{2}(z)\geq\frac{1}{2}\frac{\|z\|^{2}}{D^{2}(S)}$. Thus $\frac{1}{2}\gamma_{S}^{2}$ is $\frac{1}{D^{2}(S)}$-strongly convex w.r.t. $(0,0)$. If $\gamma_{S}(y)=0$ and $y\neq 0$, then $\gamma_{S}(ty)=t\gamma_{S}(y)=0$ for any $t>0$. Thus $\frac{1}{2}\gamma_{S}^{2}(y)$ is constant along this ray and consequently cannot be strongly convex at $y$.
@@ 3.3.2 Proof of Theorem 12
If $\gamma_{S}(y)=0$ and $y=0$, then $\partial(\frac{1}{2}\gamma_{S}^{2})(y)=\{0\}$. For any point $z$, Lemma 5 ensures $\frac{1}{2}\gamma_{S}^{2}(z)\leq\frac{1}{2}\frac{\|z\|^{2}}{R^{2}(S)}$. Thus $\frac{1}{2}\gamma_{S}^{2}$ is $\frac{1}{R^{2}(S)}$-smooth w.r.t. $(0,0)$. If $\gamma_{S}(y)=0$ and $y\neq 0$, the set $S$ is unbounded in the direction of $y$ and we still have $\partial(\frac{1}{2}\gamma_{S}^{2})(y)=\{0\}$. Since $B(0,R(S))\subseteq S$, by convexity, we have $B(ty,R(S))\subseteq S$ for any $t>0$, thus $T=\cup_{t>0}B(ty,R(S))\subseteq S$. It follows that
@@ 3.3.3 Proof of Theorem 16
Let $S$ be the convex hull of $\{0\}\cup B(c,\gamma)$. Consider $\bar{y}=(\sqrt{D^{2}-R^{2}},-R)\in\mathrm{bdry\ }S$ and $\zeta=(0,-1)\in N_{S}(\bar{y})$, then $c=\bar{y}-\gamma\zeta$. By Lemma 19, we get
@@ 3.3.4 Proof of Theorem 17
$\partial\left(\frac{1}{2}\gamma^{2}_{S}\right)(y)=\gamma_{S}(y)\partial\gamma_{S}(y)=\left\{\frac{\gamma_{S}(y)\zeta}{\zeta^{T}\bar{y}}\mid\zeta\in N_{S}(\bar{y})\right\}$
@@
$S\cap B(\bar{y},\bar{\eta})$ =$=\left\{z\mid\frac{1}{2}\gamma^{2}_{S}(z)\leq\frac{1}{2}\right\}\cap B(\bar{y},\bar{\eta})$
@@
=$=\left\{z\mid\left\|z-\bar{y}+\frac{\zeta}{\mu\zeta^{T}\bar{y}}\right\|^{2}\leq\frac{1}{\mu^{2}(\zeta^{T}\bar{y})^{2}}\right\}\cap B(\bar{y},\bar{\eta})\ .$
@@ 4 Accelerated Feasibility and Optimization Methods
where the Lipschitz constant $M$ follows from Lemma 5 and smoothness and strong convexity constants $L$ and $\mu$ for $\frac{1}{2}h_{i}^{2}$ are exactly our main results in Corollaries 13 and 14. Note the boundedness assumed of each $S_{i}$ here is not necessary for smoothness of $\gamma_{S_{i}}^{2}$ but rather a convient sufficient condition enabling use of our corollaries.
Consider an optimization problem (2) given by compact convex sets $S_{1},\dots S_{m}$ and continuous concave function $f$ with bounded level set $S_{f}=\{x\mid f(x)>0\}$. Assume $0\in\mathrm{int\ }(\cap_{i}S_{i})\cap S_{f}$ and let $G$ be the local Lipschitz constant of $f$ on $S_{f}$. Observing $f^{\Gamma}(y)=\gamma_{\mathrm{hypo\ }f}((y,1))$ where hypo $\mathrm{\text{hypo }\ }f=\{(x,u)\mid 0<u\leq f(x)\}$, the radial dual problem (7) is also minimizing a finite maximum of $n=m+1$ gauges $\max\{\gamma_{\mathrm{hypo\ }f}(y,1),\gamma_{S_{i}}(y)\}$. Supposing each $S_{i}$ is $\beta_{i}$-smooth and $\alpha_{i}$-strongly convex and $f$ is $L_{f}$-smooth and $\mu_{f}$-strongly convex on $S_{f}$222 The fact that radial methods only depend on the smoothness and strong convexity of $f$ on a level set rather than globally is convenient benefit previously noted in [2]., (7) takes the form (20) with
@@
utilizing the same constant bounds for the gauge of each $S_{i}$ as in (21) and where the Lipschitz constant for $f^{\Gamma}$ follows from [2, Proposition 1] and its strong convexity and smoothness constants follow from our Theorems 11 and 12 since Lemma 7 ensures local $\delta L_{f}$-smoothness and $\delta^{3}\mu_{f}$-strong convexity of $\mathrm{hypo\ }f$ with $1/\sqrt{1+G^{2}}\leq\delta\leq 1$.
Given these common structures, one can design optimization methods for (20), taking advantage. To the best of our knowledge, methods and theory for this exact setting have not been previously derived. Prior works have considered settings of minimizing a finite maximum of Lipschitz smooth functions [29, 30, 31, 32]. A maximum of Lipschitz strongly convex functions is itself Lipschitz and strongly convex, so guarantees like [33] apply. Below, we present three standard algorithms generalized to minimize a maximum of Lipschitz functions that, when squared, are smooth and/or strongly convex. Appendix A.3 provides convergence theorems, summarized in Table 2.
@@
Accel. Gen. Grad Mthd - - $O\left(\dfrac{LD^{2}}{p_{*}(T+1)^{2}}\right)$ $O\left((1-\sqrt{L/\mu})^{T}\right)$
@@
Consider the setting of (21). In general, applying the level projection method (25) with $\bar{h}=1$ converges to feasibility with
@@
and if $\alpha>0$ and $T+1\geq\frac{32}{R^{2}\mu}\log(\mu\ \mathrm{dist}(y_{0},\cap S_{i})^{2})$, this improves to
@@
$\max_{i}\gamma_{S_{i},e_{i}}(y_{T})-h(y^{*})\leq\frac{h(y_{0})^{2}-h(y^{*})^{2}+\frac{t_{0}^{2}L}{1-t_{0}}\|y_{0}-y^{*}\|^{2}}{2h(y^{*})}\min\left\{\left(1-\sqrt{\frac{\mu}{L}}\right)^{T},\ \frac{4}{\left(2+T\sqrt{\frac{t_{0}^{2}}{1-t_{0}}}\right)^{2}}\right\}.$
@@
Consider the setting of (22) and let $x^{*}$ maximize (2). Let $x_{0}=y_{0}=0$, $y_{k}$ denote the iterates of some method minimizing the radial dual problem, and $x_{k}=y_{k}/h(y_{k})\in\cap S_{i}$. In general, the iterates $y_{k}$ of the subgradient method (23) with $s_{k}=\frac{D}{\|h(y_{k})g_{k}\|\sqrt{T+1}}$ converges to optimality with
@@
and if $\alpha>0$ and $s_{k}=\frac{2}{\mu(k+2)+\frac{M^{4}}{\mu(k+1)}}$, this improves to
@@
$\frac{f(x^{*})-f(x_{T})}{f(x^{*})}\leq\frac{1}{2}\left(\frac{f(x^{*})^{2}}{f(x_{0})^{2}}-1+\frac{t_{0}^{2}L}{1-t_{0}}\|x_{0}-x^{*}\|^{2}\right)\min\left\{\left(1-\sqrt{\frac{\mu}{L}}\right)^{T},\ \frac{4}{\left(2+T\sqrt{\frac{t_{0}^{2}}{1-t_{0}}}\right)^{2}}\right\}\ .$
@@
$h(y_{T})-h(y^{*})\leq\frac{h(y_{0})^{2}-h(y^{*})^{2}+\frac{t_{0}^{2}L}{1-t_{0}}\|y^{*}\|^{2}}{2h(y^{*})}\min\left\{\left(1-\sqrt{\frac{\mu}{L}}\right)^{T},\ \frac{4}{\left(2+T\sqrt{\frac{t_{0}^{2}}{1-t_{0}}}\right)^{2}}\right\}\ .$
To translate these radial dual optimality guarantees to the original (primal) problem, observe that $x_{k}=y_{k}/h(y_{k})$ has $f(x_{k})\geq 1/f^{\Gamma}(y_{k})$ and $x_{k}\in\cap S_{i}$ and $x^{*}$ has $f(x^{*})=1/h(y^{*})$. Hence each result above upper bounds $1/f(x_{k})-1/f(x^{*})$ by some $\epsilon$. The claims then follow as this implies $(f(x^{*})-f(x_{k}))/f(x^{*})\leq f(x^{*})\epsilon$. ∎
@@
Further Methods and Oracle Models (When $n$ is Large). Lastly, since the level projection and generalized gradient methods scale poorly with $n$ (i.e., requiring QP solves), we note two methods applicable when $n$ is large and each $h_{i}$ when squared is $L$-smooth. The accelerated smoothing method of [29, 30] replaces the finite maximum by a $1/\epsilon$-smooth approximation. Then applying Nesterov’s accelerated method gives an overall convergence rate of $O(L/T)$, improving on the subgradient method’s $O(1/\sqrt{T})$. Renegar and Grimmer [34]’s restarting scheme showed such a method can attain a faster $L/\mu\log(1/\epsilon)$, when $\mu$-strong convexity additionally holds. Alternatively, replacing the finite maximum by $\max_{\theta\in\Delta}\sum\theta_{i}h_{i}(x)$ where $\Delta$ is the simplex gives an equivalent convex-concave minimax optimization problem. Then the same $O(L/T)$ speed up can be attained by the extragradient method [31].

### v4

@@ 3.1 Examples
$\mu$ $=\inf_{\|\bar{y}-b\|_{2}=1}\frac{1}{(\bar{y}-b)^{T}\bar{y}+\|\bar{y}\|^{2}_{2}}=\inf_{\|\bar{y}-b\|_{2}=1}\frac{1}{2+3(\bar{y}-b)^{T}b+\|b\|^{2}_{2}}=\frac{1}{(1+\|b\|_{2})(2+\|b\|_{2})}\ ,$
$L$ $=\sup_{\|\bar{y}-b\|_{2}=1}\frac{(\bar{y}-b)^{T}\bar{y}+\|\bar{y}\|_{2}^{2}}{((\bar{y}-b)^{T}\bar{y})^{3}}=\sup_{\|\bar{y}-b\|_{2}=1}\frac{2+3(\bar{y}-b)^{T}b+\|b\|^{2}_{2}}{(1+(\bar{y}-b)^{T}b)^{3}}=\frac{2-\|b\|_{2}}{(1-\|b\|_{2})^{2}}\ ,$
where the infimum and supremum are taken by Cauchy-Schwarz $\|b\|_{2}\leq(\bar{y}-b)^{T}b\leq\|b\|_{2}$. Then following Lemma 15, $\frac{1}{2}\gamma_{E_{2}}^{2}$ is $\frac{\lambda_{\min}(A^{T}A)}{(1+\|b\|_{2})(2+\|b\|_{2})}$-strongly convex and $\frac{\lambda_{\max}(A^{T}A)(2-\|b\|_{2})}{(1-\|b\|_{2})^{2}}$-smooth.
@@ 3.3 Proofs of Theorems Characterizing Structured Gauges
Both strongly convex sets and smooth sets are defined in terms of balls built by a boundary point and corresponding normal vector. This perspective is critical for the proofs of our main theorems. The following lemmas characterize the gauges of these approximating balls. Their proofs are deferred to Appendix A.2.5 and A.2.6. Note that below, we consider sets $B(\bar{y}-r\zeta,r)$, which may not contain the origin. As a result, it is important that we defined the gauge as $\inf\{\lambda>0\mid x\in\lambda S\}$, which may no longer equal $\sup\{\lambda>0\mid x\not\in\lambda S\}$. Note when $0\notin S$, $\gamma_{S}(0)=\infty$.
@@ 3.3.1 Proof of Theorem 11
$\partial\left(\frac{1}{2}\gamma^{2}_{B}\right)(y)=\gamma_{B}(y)\partial\gamma_{B}(y)=\left\{\frac{\gamma_{B}(y)\zeta}{\zeta^{T}\bar{y}}\right\}\ .$ (21)
@@
Since $S$ is strongly convex w.r.t. $(\bar{y},\zeta)$, then $S\cap B(\bar{y},\bar{\eta})\subseteq B\left(\bar{y}-\frac{1}{\alpha}\zeta,\frac{1}{\alpha}\right)\cap B(\bar{y},\bar{\eta})$ for some $\bar{\eta}$. Hence for $z$ with $\|z-y\|\leq\gamma_{S}(y)\frac{R(S)}{D(S)+R(S)}\bar{\eta}$, the boundary point $\bar{z}=z/\gamma_{S}(z)$ has
$\|\bar{z}-\bar{y}\|$ $=\|\frac{z}{\gamma_{S}(z)}-\frac{y}{\gamma_{S}(y)}\|$
$\leq\frac{\|z\||\gamma_{S}(y)-\gamma_{S}(z)|}{\gamma_{S}(z)\gamma_{S}(y)}+\frac{\|z-y\|}{\gamma_{S}(y)}$
$\leq\frac{D(S)\|z-y\|}{R(S)\gamma_{S}(y)}+\frac{\|z-y\|}{\gamma_{S}(y)}$
$\leq\bar{\eta}\ ,$
where the second inequality is given by $\gamma_{S}$ being $1/R(S)$-Lipschitz and $\|\bar{z}\|\leq D(S)$. Then $\bar{z}\in S\cap B(\bar{y},\bar{\eta})$ implies $\bar{z}\in B\left(\bar{y}-\frac{1}{\alpha}\zeta,\frac{1}{\alpha}\right)$, it follows $\gamma_{S}(z)\geq\gamma_{B}(z)$. Applying this bound and Lemma 19 (a) and (b), we conclude for any $\tilde{\mu}<\mu$ in some neighborhood of $y$, all $z$ have
$\frac{1}{2}\gamma_{S}^{2}(z)\geq\frac{1}{2}(\gamma^{B})^{2}(z)\geq\frac{1}{2}\gamma_{S}^{2}(y)+g^{T}(z-y)+\frac{\tilde{\mu}}{2}\|z-y\|^{2}\ .$ (22)
@@
If $\gamma_{S}(y)=0$ and $y=0$, then $\partial(\frac{1}{2}\gamma_{S}^{2})(y)=\{0\}$. For any point $z$, Lemma 5 ensures $\frac{1}{2}\gamma_{S}^{2}(z)\geq\frac{1}{2}\frac{\|z\|^{2}}{D^{2}(S)}$. Thus $\frac{1}{2}\gamma_{S}^{2}$ is $\frac{1}{D^{2}(S)}$-strongly convex w.r.t. $(0,0)$.
If $\gamma_{S}(y)=0$ and $y\neq 0$, then $\gamma_{S}(ty)=t\gamma_{S}(y)=0$ for any $t>0$. Thus $\frac{1}{2}\gamma_{S}^{2}(y)$ is constant along this ray and consequently cannot be strongly convex at $y$.
@@ 3.3.2 Proof of Theorem 12
If $\gamma_{S}(y)=0$ and $y=0$, then $\partial(\frac{1}{2}\gamma_{S}^{2})(y)=\{0\}$. For any point $z$, Lemma 5 ensures $\frac{1}{2}\gamma_{S}^{2}(z)\leq\frac{1}{2}\frac{\|z\|^{2}}{R^{2}(S)}$. Thus $\frac{1}{2}\gamma_{S}^{2}$ is $\frac{1}{R^{2}(S)}$-smooth w.r.t. $(0,0)$.
If $\gamma_{S}(y)=0$ and $y\neq 0$, the set $S$ is unbounded in the direction of $y$ and we still have $\partial(\frac{1}{2}\gamma_{S}^{2})(y)=\{0\}$. Since $B(0,R(S))\subseteq S$, by convexity, we have $B(ty,R(S))\subseteq S$ for any $t>0$, thus $T=\cup_{t\geq 0}B(ty,R(S))\subseteq S$. Since $T$ is unbounded in the direction of $y$, and bounded in the direction of y-y, for $\bar{z}=z/\gamma_{T}(z)$, we have the following two cases. If $y^{T}z\geq 0$, then $\gamma_{T}(z)=\frac{\|z\|}{\|\bar{z}\|}=\frac{1}{R(S)}\|z-\frac{y^{T}z}{\|y\|^{2}}y\|$. If $y^{T}z<0$, then $\gamma_{T}(z)=\gamma_{B(0,R(S))}(z)=\frac{\|z\|}{R(S)}$. It follows that
@@ 3.3.3 Proof of Theorem 16
Let $S$ be the convex hull of $B(0,\frac{\zeta^{T}\bar{y}}{2})\cup B(c,1/\gamma)$. Consider $\bar{y}=(\sqrt{D^{2}-R^{2}},-R)\in\mathrm{bdry\ }S$ and $\zeta=(0,-1)\in N_{S}(\bar{y})$, then $c=\bar{y}-\zeta/\gamma$. Note that since $B(0,\frac{\zeta^{T}\bar{y}}{2})$ has radius $\frac{\zeta^{T}\bar{y}}{2}<\zeta^{T}\bar{y}$, one has $(0,-1)\in N_{S}(\bar{y})$. By Lemma 19, we get
@@ 3.3.4 Proof of Theorem 17
$\partial\left(\frac{1}{2}\gamma^{2}_{S}\right)(y)=\gamma_{S}(y)\partial\gamma_{S}(y)=\left\{\frac{\gamma_{S}(y)\zeta}{\zeta^{T}\bar{y}}\mid\zeta\in N_{S}(\bar{y})\right\}$ (23)
@@
$S\cap B(\bar{y},\bar{\eta})$ $=\left\{z\mid\frac{1}{2}\gamma^{2}_{S}(z)\leq\frac{1}{2}\right\}\cap B(\bar{y},\bar{\eta})$
@@
$=\left\{z\mid\left\|z-\bar{y}+\frac{\zeta}{\mu\zeta^{T}\bar{y}}\right\|^{2}\leq\frac{1}{\mu^{2}(\zeta^{T}\bar{y})^{2}}\right\}\cap B(\bar{y},\bar{\eta})\ .$
@@ 4 Accelerated Feasibility and Optimization Methods
where the Lipschitz constant $M$ follows from Lemma 5 and smoothness and strong convexity constants $L$ and $\mu$ for $\frac{1}{2}h_{i}^{2}$ are exactly our main results in Corollaries 13 and 14. Note the boundedness assumed of each $S_{i}$ here is not necessary for smoothness of $\gamma_{S_{i}}^{2}$ but rather a convenient sufficient condition enabling use of our corollaries.
Consider an optimization problem (2) given by compact convex sets $S_{1},\dots S_{m}$ and continuous concave function $f$ with bounded level set $S_{f}=\{x\mid f(x)>0\}$. Assume $0\in\mathrm{int\ }(\cap_{i}S_{i})\cap S_{f}$ and let $G$ be the local Lipschitz constant of $f$ on $S_{f}$. Observing $f^{\Gamma}(y)=f_{\Gamma}(y)$ [1, Proposition 11] and $f_{\Gamma}(y)=\inf\{v>0\mid f(y/v)>1/v\}=\inf\{v>0\mid\frac{(y,1)}{v}\in\mathrm{hypo\ }f\}=\gamma_{\mathrm{hypo\ }f}((y,1)),$ where $\mathrm{\text{hypo }\ }f=\{(x,u)\mid 0<u\leq f(x)\}$, the radial dual problem (7) is also minimizing a finite maximum of $n=m+1$ gauges $\max\{\gamma_{\mathrm{hypo\ }f}(y,1),\gamma_{S_{i}}(y)\}$. Supposing each $S_{i}$ is $\beta_{i}$-smooth and $\alpha_{i}$-strongly convex and $f$ is $L_{f}$-smooth and $\mu_{f}$-strongly convex on $S_{f}$ [^2] The fact that radial methods only depend on the smoothness and strong convexity of $f$ on a level set rather than globally is convenient benefit previously noted in [2]., (7) takes the form (24) with
@@
utilizing the same constant bounds for the gauge of each $S_{i}$ as in (25) and where the Lipschitz constant for $f^{\Gamma}$ follows from [2, Proposition 1] and its strong convexity and smoothness constants follow from our Theorems 11 and 12 since Lemma 7 ensures local $|\delta|L_{f}$-smoothness and $|\delta|^{3}\mu_{f}$-strong convexity of $\mathrm{hypo\ }f$ with $1/\sqrt{1+G^{2}}\leq|\delta|\leq 1$.
One can design optimization methods for (24), taking advantage of these basic structures. To the best of our knowledge, methods and theory for this exact setting have not been previously derived. Prior works have considered settings of minimizing a finite maximum of Lipschitz smooth functions [29, 30, 31, 32]. A maximum of Lipschitz strongly convex functions is itself Lipschitz and strongly convex, so guarantees like [33] apply. Below, we present three standard algorithms generalized to minimize a maximum of Lipschitz functions that, when squared, are smooth and/or strongly convex. Appendix A.3 provides convergence theorems, summarized in Table 2.
@@
Accel. Gen. Grad Mthd - - $O\left(\dfrac{LD^{2}}{p_{*}(T+1)^{2}}\right)$ $O\left((1-\sqrt{\mu/L})^{T}\right)$
@@
Consider the setting of (25), and let $R=\min\{R(S_{i}-e_{i})\}$. In general, applying the level projection method (29) with $\bar{h}=1$ converges to feasibility with
@@
and if $\mu>0$ and $T+1\geq\frac{32}{R^{2}\mu}\log(\mu\ \mathrm{dist}(y_{0},\cap S_{i})^{2})$, this improves to
@@
$\max_{i}\gamma_{S_{i},e_{i}}(y_{T})-h(y^{*})\leq\frac{h(y_{0})^{2}-h(y^{*})^{2}+\gamma_{0}\|y_{0}-y^{*}\|^{2}}{2h(y^{*})}\min\left\{\left(1-\sqrt{\frac{\mu}{L}}\right)^{T},\ \frac{4L}{\left(2\sqrt{L}+T\sqrt{\gamma_{0}}\right)^{2}}\right\},$
where $\gamma_{0}=\frac{t_{0}(t_{0}L-\mu)}{1-t_{0}}$.
@@
Consider the setting of (26), let $x^{*}$ maximize (2), and $R=\min\{R(S_{i}),R(S_{f})\}$. Let $x_{0}=y_{0}=0$, $y_{k}$ denote the iterates of some method minimizing the radial dual problem, and $x_{k}=y_{k}/h(y_{k})\in\cap S_{i}$. In general, the iterates $y_{k}$ of the subgradient method (27) with $s_{k}=\frac{D}{\|h(y_{k})g_{k}\|\sqrt{T+1}}$ converges to optimality with
@@
and if $\alpha>0$ and $s_{k}=\frac{2}{\mu(k+2)+\frac{1}{R^{4}\mu(k+1)}}$, this improves to
@@
$\frac{f(x^{*})-f(x_{T})}{f(x^{*})}\leq\frac{1}{2}\left(\frac{f(x^{*})^{2}}{f(x_{0})^{2}}-1+\gamma_{0}\|x_{0}-x^{*}\|^{2}\right)\min\left\{\left(1-\sqrt{\frac{\mu}{L}}\right)^{T},\ \frac{4L}{\left(2\sqrt{L}+T\sqrt{\gamma_{0}}\right)^{2}}\right\}\ ,$
where $\gamma_{0}=\frac{t_{0}(t_{0}L-\mu)}{1-t_{0}}$.
@@
$h(y_{T})-h(y^{*})\leq\frac{h(y_{0})^{2}-h(y^{*})^{2}+\gamma_{0}\|y^{*}\|^{2}}{2h(y^{*})}\min\left\{\left(1-\sqrt{\frac{\mu}{L}}\right)^{T},\ \frac{4L}{\left(2\sqrt{L}+T\sqrt{\gamma_{0}}\right)^{2}}\right\}\ .$
To translate these radial dual optimality guarantees to the original (primal) problem, observe that $x_{k}=y_{k}/h(y_{k})$ and $h(y_{k})\geq f^{\Gamma}(y)=\sup\{v>0\mid v\cdot f(y_{k}/v)\leq 1\}$ have $f(x_{k})\geq 1/h(y_{k})$ and $x_{k}\in\cap S_{i}$ and $x^{*}$ has $f(x^{*})=1/h(y^{*})$. Hence each result above upper bounds $1/f(x_{k})-1/f(x^{*})$ by some $\epsilon$. The claims then follow as this implies $(f(x^{*})-f(x_{k}))/f(x^{*})\leq f(x_{k})\epsilon\leq f(x^{*})\epsilon$. ∎
@@
Further Methods and Oracle Models (When $n$ is Large). Lastly, since the level projection and generalized gradient methods scale poorly with $n$ (i.e., requiring QP solves), we note two methods applicable when $n$ is large and each $h_{i}$ when squared is $L$-smooth. The accelerated smoothing method of [29, 30] replaces the finite maximum by a $1/\epsilon$-smooth approximation. Then applying Nesterov’s accelerated method gives an overall convergence rate of $O(L/T)$, improving on the subgradient method’s $O(1/\sqrt{T})$. Renegar and Grimmer [34]’s restarting scheme showed such a method can attain a faster $O(1/\sqrt{\epsilon})$ rate, when $\mu$-strong convexity additionally holds. Alternatively, replacing the finite maximum by $\max_{\theta\in\Delta}\sum\theta_{i}h_{i}(x)$ where $\Delta$ is the simplex gives an equivalent convex-concave minimax optimization problem. Then the same $O(L/T)$ speed up can be attained by the extragradient method [31].